OpenStudy (anonymous):

find x for 6x-5, 2x+35

OpenStudy (asadkarim7):

is 6x-5 = 2x+35??

OpenStudy (anonymous):

i think the problem is wrong written

OpenStudy (anonymous):

its an image here

OpenStudy (asadkarim7):

hey its vertically opposite angles

OpenStudy (asadkarim7):

so 6x-5 =2x+35 4x=40 x=10 degrees

OpenStudy (anonymous):

how did you get 40?

OpenStudy (asadkarim7):

shift -5 to other side , it would be +5 35+5=40

OpenStudy (anonymous):

Here is another problem also

OpenStudy (asadkarim7):

they are also equal coz they are corrorponding angles

OpenStudy (asadkarim7):

corresponding angles are also equal so 4x-8= 2x+16 4x-2x=16+8 2x=24 x=12

OpenStudy (anonymous):

thnk you

OpenStudy (anonymous):

how many degrees from the horizontal position are the wings of the airplane

OpenStudy (asadkarim7):

the wings are 63deg from vertical so it would be 90-63 =27 deg from horizontal

OpenStudy (anonymous):

your great thnx lol